You can afford monthly deposits of $210 into an account that pays 2.7% compounded monthly. How long until you have $8,500 to buy a boat?

Type the number of months: ___

(Please square, label, or circle the answer)
Answer #1

Correct answer will be 138.91 Month

Detail working for your refrence

A= P(1+r)^t
8500=210(1+0.027)^t
8500/210= (1.027)^t
40.476=(1.027)^t
take log on both side
40.476=(1.027)^t
log40.476=tlog1.027
1.60720=tX0.01157
t=1.60720/0.01157=138.91 Month